It takes 585 J of energy to raise the temperature of 125.6 g mercury from 20.08C to 53.58C. Calculate the specific heat capacity and the molar heat capacity of mercury.
1
Expert's answer
2020-12-01T09:31:59-0500

The specific heat capacity is the amount of energy (in form of heat) that must be added to the unit of mass in order to rise its temperature by 1 degree:

"c = \frac{Q}{m∆T} = \frac{585\text{ J}}{125.6\text{ g}·(53.58-20.08)\text{ K}} = 0.139" J g-1 K-1.

The molar heat capacity , in its turn, is the amount of energy (in form of heat) that must be added to 1 mole of the substance in order to rise its temperature by 1 degree.

The number of the moles of mercury in 125.6 g is (molar mass of mercury is 200.59 g/mol):

"n = \frac{m}{M} = \frac{125.6\text{ g}}{200.59\text{ g/mol}} = 0.626" mol.

Then, the molar heat capacity of mercury is:

"c_m = \frac{Q}{n∆T} = \frac{585\text{ J}}{0.626 \text{ mol}·(53.58-20.08)\text{ K}} = 27.89" J mol-1 K-1.

Answer: the specific heat capacity and the molar heat capacity of mercury are 0.139 J g-1 K-1 and 27.89 J mol-1 K-1, respectively.


Note: remember that the difference temperature in celsius and in kelvin have the same numeric value due to the conversion rule (+273.15 to convert from celsius to kelvin).